OpenStudy (msmr):

Use algebra to find each point at which the line (x-2y=-5) intersects the circle (x^2 + y^2 = 20)

OpenStudy (msmr):

I know the answer is (3,4) and (-5,0), but I have no idea how to get there.

OpenStudy (anonymous):

\[x-2y=-5\] \[x=2y-5\] replace in second equation get \[(2y-5)^2+y^2=20\] solve this quadratic equation in y

OpenStudy (anonymous):

some mistake in your answer or your question. i am going to guess that the circle should be \[x^2+y^2=25\]am i right?

OpenStudy (msmr):

oh, you're right, it is x^2 + y^2 = 25, sorry!

OpenStudy (anonymous):

because that would give you \[(2y-5)^2+y^2=25\] and the solutions would be \[y=0, y=4\] and in fact (3,4) is not on the circle you wrote

OpenStudy (msmr):

The back of my math book says the answer is (3,4),(-5,0)...so I'm not sure...

OpenStudy (anonymous):

yeah that is fine

OpenStudy (anonymous):

solve the quadratic equation \[(2y-5)^2+y^2=25\] get \[5y^2-20y+25=25\] \[5y^2-20y=0\] \[5y(y-4)=0\] \[y=0, y =4\]

OpenStudy (msmr):

and then how do you find the x coordinate? do you plug it back in to the line equation?

OpenStudy (anonymous):

replace y by 0 in the line, get x = -5, replace y by 4 in the line, get x = 3

OpenStudy (anonymous):

right, back in to the line equation

OpenStudy (msmr):

thank you so much!

OpenStudy (anonymous):

yw